Given the following circuit, find the voltage V3, of the circuit, if element A is

10 Ω 20 Ω 20 V S 60V A 1, 40 Ω + 30 Ω

a) A short circuit

b) An independent 9V voltage source, with the positive reference on the left

c) A dependent current source, with the arrowhead to the left and marked 5i1.
Answer #1

sol til 20v gou IA 60V itiz 2 402 + 3or VA go 202 = IA 20 (eti,)-(1+iz) current through the 4or = ع -6 4on 2012 - 40(1-12) -b nou 901 20V lusti 1200) av f 60v Liz+190n) 12 nou 30 VB -20 –9 +12on120)=0 ison 29 a 1.45A. 20 Чол - 1.45 +i, -liq+1.45) iilor e) 20V 20~ Siniz 51,-iq+i, 6i-12 t sii Gor ya -ie Wor 901 N3 KULA 20(51,-12) --20 +40 (1,-12)-3012 = 0 looi 2012 +401,-4